OpenStudy (anonymous):

if x=-1 then what is y. y=-x^2-5x-4

OpenStudy (anonymous):

plug and chug y=-(-1)^2-5(-1)-4

OpenStudy (anonymous):

what do the 2 negatives mean

OpenStudy (anonymous):

y²=(-1)²+5-4=1+5-4=2<=> y²=2 implies that y=-2 or y=2

OpenStudy (anonymous):

sry y=sqrt(2) or -sqrt(2)

OpenStudy (anonymous):

so its 2 or -2

OpenStudy (anonymous):

no sqrt(2) or -sqrt(2)

OpenStudy (anonymous):

its not sqrt

OpenStudy (anonymous):

linear to quadratic

OpenStudy (anonymous):

y=2,-2

OpenStudy (anonymous):

y.y or just y ?

OpenStudy (anonymous):

just y

OpenStudy (anonymous):

ok then y=2 or y=-2

OpenStudy (anonymous):

its a chart so x=-1

OpenStudy (anonymous):

you plug it in and somehow get y

OpenStudy (anonymous):

can anybody help me

OpenStudy (anonymous):

y=0